### v2

# Steering Dialogue Dynamics for Robustness against Multi-turn Jailbreaking Attacks

Abstract
Large language models (LLMs) are shown to be vulnerable to jailbreaking attacks where adversarial prompts are designed to elicit harmful responses. While existing defenses effectively mitigate single-turn attacks by detecting and filtering unsafe inputs, they fail against multi-turn jailbreaks that exploit contextual drift over multiple interactions, gradually leading LLMs away from safe behavior. To address this challenge, we propose a safety steering framework grounded in safe control theory, ensuring invariant safety in multi-turn dialogues. Our approach models the dialogue with LLMs using state-space representations and introduces a novel neural barrier function (NBF) to detect and filter harmful queries emerging from evolving contexts proactively. Our method achieves invariant safety at each turn of dialogue by learning a safety predictor that accounts for adversarial queries, preventing potential context drift toward jailbreaks. Extensive experiments under multiple LLMs show that our NBF-based safety steering outperforms safety alignment, prompt-based steering and lightweight LLM guardrails baselines, offering stronger defenses against multi-turn jailbreaks while maintaining a better trade-off among safety, helpfulness and over-refusal. Check out the website here https://sites.google.com/view/llm-nbf/home. Our code is available on https://github.com/HanjiangHu/NBF-LLM.
Warning: This paper contains examples of harmful LLM responses.
1 Introduction
Figure 1: Single-turn vs multi-turn jailbreaks.
Despite the tremendous potential of large language models (LLMs) across a variety of applications, frontier models remain vulnerable to jailbreaking attacks, wherein adversarial prompts are designed to elicit harmful responses (wei2023jailbroken, ; anwar2024foundational, ; sun2024trustllm, ). These attacks include optimization-based methods (zou2023universal, ; geisler2024attacking, ; andriushchenko2024jailbreaking, ) and automated techniques in which attackers use LLMs to produce jailbreaks (chao2023jailbreaking, ; liu2024autodan, ; mehrotra2023tree, ). To counter these threats, defenses such as fine-tuning-based algorithms (yuan2024refuse, ; zou2024improving, ), inference-time interventions (arditi2024refusal, ; bhattacharjee2024towards, ; robey2023smoothllm, ; li2024inference, ), and reasoning-based guardrails (kang2024r, ; liu2025guardreasoner, ; zaremba2025trading, ) have been proposed. These defenses have been successful in reducing the effectiveness of single-turn attacks, wherein an adversary can only attempt to jailbreak a targeted model via a single round of conversation.
Figure 2: Overview of safety steering based on neural dialogue dynamics and barrier function.
Unfortunately, the success of current defenses against single-turn jailbreaks has not extended to the more sophisticated setting of multi-turn jailbreaking, wherein an attacker attempts to elicit harmful content throughout multiple rounds of conversation (li2024llm, ; pavlova2024automated, ). In the multi-turn setting, attackers exploit the gradual shift in context to bypass safeguards (ren2024derail, ; russinovich2024great, ; jiang2024red, ), making detection and mitigation significantly more challenging, as shown in Figure˜1. Multi-turn jailbreaks use adaptive and dynamic interactions to subtly steer LLMs toward unsafe outputs (zhou2024speak, ; liu2024imposter, ), circumventing traditional single-turn defenses. One concurrent defense method (lu2025x, ) proposes to learn more fine-grained boundary-safe and harmful representations through safety fine-tuning, but the multi-turn dialogue context is not explicitly considered and the fine-tuning method is not agnostic to different LLMs as a versatile guardrail (markov2023holistic, ; zeng2024shieldgemma, ; inan2023llama, ).
In this work, we approach the problem of defending against multi-turn jailbreaks through the lens of dynamical systems. We view the context of each dialogue turn as an evolving hidden state and define a successful attack as one that transitions into an unsafe region of the state space. Inspired by forward invariance in safe control theory (liu2014control, ; ames2019control, ; robey2020learning, ), we introduce the concept of invariant safety for dialogue dynamics, ensuring that at every step of a conversation, the system remains within a user-specified safe set, thereby preventing attackers from gradually leading LLMs to generate harmful responses. As shown in Figure˜2, we first learn state-space representations corresponding to the neural dialogue dynamics from multi-turn conversations. Then, given these representations, we train a safety predictor as a neural barrier function (NBF) to predict whether the states corresponding to a multi-turn conversation drift into harmful regions of the state space. During the evaluation phase, the NBF filters out potentially harmful queries based on the predictor’s outputs. Extensive experiments validate the effectiveness of the proposed NBF-based steering and show strong generalizability to different LLMs and multi-turn jailbreaking attacks. In summary, the contributions are listed below.
  • We proposed a control-theoretical framework to model the neural dialogue dynamics with LLMs and achieve invariant safety against multi-turn jailbreaking attacks.
  • We introduce a neural barrier function (NBF) that evaluates the potential safety violation given the worst-case harmful query within the current dialogue context at each turn.
  • Comprehensive experiments show that the proposed NBF-based safety steering can outperform defense baselines with a better trade-off of safety, helpfulness and over-refusal on multiple LLMs.
2 Related Work
LLM Jailbreaking Attacks and Defenses.
Jailbreak attacks on LLMs have advanced from hand-crafted prompts to automated red-teaming approaches. Optimization-based methods, including gradient-based and evolutionary attacks (zou2023universal, ; geisler2024attacking, ; liu2023autodan, ; andriushchenko2024jailbreaking, ), generate adversarial inputs, while automated attackers leverage LLMs for iterative refinements (chao2023jailbreaking, ; liu2024autodan, ; robey2024jailbreaking, ). While existing safety alignment (yuan2024refuse, ; zou2024improving, ; zhang2025safety, ), inference-time steering methods (arditi2024refusal, ; bhattacharjee2024towards, ) and reasoning-based LLM guardrails (kang2024r, ; liu2025guardreasoner, ) are effective against various single-turn jailbreaks, they struggle to defend against multi-turn jailbreaks (wang2024mrj, ; tong2024securing, ). Multi-turn jailbreaking scenarios are developed by embedding malicious intent gradually (jiang2024red, ), breaking down harmful prompts into benign sub-queries (yu2024cosafe, ; zhou2024speak, ; liu2024imposter, ), designing attack patterns (ren2024derail, ), and dynamically adjusting attack queries based on contextual feedback (li2024llm, ; yang2024chain, ; russinovich2024great, ). One contemporaneous multi-turn defense method (lu2025x, ) learns the safety boundary through fine-tuning without explicitly considering multi-turn context, and also cannot be used as a guard model across different LLMs. Current guard models typically introduce a separate model designed to moderate LLMs to filter out unsafe content (markov2023holistic, ; zeng2024shieldgemma, ; inan2023llama, ), but dynamically drifting context poses a challenge to existing reactive safety defense mechanisms. To the best of our knowledge, we are the first to safeguard LLM dialogues dynamically from jailbreaks through online filtering of harmful prompts before the prompts are sent to LLMs.
Learning-based Safe Control with Neural Certificates.
In the literature of control and robotics, there is extensive research on learning-based controllers for dynamical systems that provide safety guarantees or certificates (boffi2021learning, ; herbert2021scalable, ; xiao2023barriernet, ; lindemann2021learning, ; chang2019neural, ; mazouz2022safety, ). Neural networks have been employed to parameterize control barrier functions (CBFs) to achieve forward invariance (robey2020learning, ; so2023train, ; zinage2023neural, ; dawson2022safe, ; dai2022learning, ): Once the system states enter the user-defined safe set, they remain within it indefinitely, thereby guaranteeing safety with neural certificates. Although neural CBFs can be successfully learned and verified for control-affined dynamical systems (chen2024learning, ; hu2024verification, ; mathiesen2022safety, ; wang2023simultaneous, ; rickard2025data, ), it is still challenging to guarantee the safety for non-analytical dynamical systems in latent space (hu2024real, ; wei2022safe, ; liu2023model, ; li2023system, ; shen2024bab, ; cheng2024robust, ), as the problem we tackle here. Recently, bajcsy2024human formulate a human-AI safety filter framework and miyaoka2024cbf introduce CBFs for LLM safety at the token level against single-turn jailbreaks. However, no existing work has explored dialogue-level safety for LLMs from the perspective of neural CBFs.
3 Problem Formulation
In this section, we formalize our approach to defending against multi-turn jailbreaks by enforcing invariant safety in the conversation with LLMs. Our framework models the conversation as an evolving dynamical system, where the hidden state represents the drifting context (e.g. the production process in Figure˜1), and each dialogue turn represents a state transition influenced by user queries and LLM responses in the language space. Given the language space $\mathcal{S}$}, at each turn $k$, LLMs receive a query $U_{k}\in{\mathcal{S}}$ from users and make a response $Z_{k}\in{\mathcal{S}}$ to that query, followed by the next round of user query $U_{k+1}\in{\mathcal{S}}$ and LLMs response $Z_{k+1}\in{\mathcal{S}}$. In accordance with the AI usage policies (openai2022usage, ), LLM responses $Z_{k},k-1,2,\dots,K$ should follow the AI safety rules and fall into the safe region specified by the user ${\mathcal{S}}_{0}$. However, in the multi-turn setting, the context drift of the query along the dynamic conversation may increase the vulnerability of the LLMs for jailbreaks, even though the response is safe at each turn. To this end, we introduce the concept invariant safety and associated measures to guarantee that once the LLM response falls into a safety region – which needs to be computed, the following LLM responses will stay within it no matter what future queries are given along the query context flow.
Definition 3.1 (Invariant Safety in Multi-turn Conversation).
Given a trajectory of user queries $U_{k}\in{\mathcal{S}}$ and LLM responses $Z_{k}\in{\mathcal{S}},k=1,2,\dots,K$ and a user-specified safety region ${\mathcal{S}}_{0}\subset{\mathcal{S}}$, the query context set ${\mathcal{S}}_{context}^{(k)}$ is defined as all reasonable queries at turn $k+1$ based on previous conversation context by turn $k$, drifting from random initial context ${\mathcal{S}}_{context}^{(0)}$. The LLM is invariantly safe (i.e., will not be jailbroken in drifting context) if there exists a safety invariance set ${\mathcal{S}}_{I}\subset{\mathcal{S}}_{0}$ such that the following holds,
$\forall k=1,2,\dots,K,\forall Z_{1},\dots,Z_{k}\in{\mathcal{S}}_{I}\Rightarrow Z_{k+1}\in{\mathcal{S}}_{I},\forall U_{k+1}\in{\mathcal{S}}_{context}^{(k)}.$ (1)
For any safe but non-invariant responses $Z^{\prime}_{k}\in{\mathcal{S}}_{0}\setminus{\mathcal{S}}_{I}$ (e.g. the LLM response at turn 3 in Figure˜1), there exists a potentially harmful query $U^{\prime}_{k+1}\in{\mathcal{S}}_{context}^{(k)}$ such that the next LLM response $Z^{\prime}_{k+1}$ will inevitably go out of ${\mathcal{S}}_{0}$, resulting in LLM jailbreaks. Therefore, the safety invariance subset ${\mathcal{S}}_{I}$ is introduced to avoid non-invariant responses to achieve invariant safety against multi-turn jailbreaking attacks. In the following section, we first model the neural dialogue dynamics of multi-turn conversation with LLMs using the state-space representations. Then, we introduce the safety predictor based neural barrier function, followed by an invariant safety certificate learning framework.
Figure 3: Conversation in the language space and state-space representations in the hidden state and embedding space.
4 Methodology
4.1 Neural Dialogue Dynamics of Multi-turn Conversation with LLMs
Different from studying large language models (LLMs) dynamics of token generation in the literature (soatto2023taming, ; liu2024meanings, ; kong2024aligning, ; miyaoka2024cbf, ), we focus on the multi-turn human-AI interactive sentence-wise dynamics in the sentence embedding space and formulate it based on the state space representations in control theory. Given a $K$-turn conversation dialog with user query sentences $U_{k}\in{\mathcal{S}},k=1,\dots K,$ and response sentences of LLMs $Z_{k}\in{\mathcal{S}},k=1,\dots K,$ we first map them from the sentence language space $\mathcal{S}$} to the semantic and meaningful embedding space $\mathbb{R}^{n}$ using the pretrained sentence embedding model $f_{embedding}:{\mathcal{S}}\rightarrow\mathbb{R}^{n}$ (reimers2019sentence, ; fonseca2025safeguarding, ) as follows,
$u_{k}\in\mathbb{R}^{n}=f_{embedding}(U_{k}),z_{k}\in\mathbb{R}^{n}=f_{embedding}(Z_{k}),k=1,2,\dots,K.$ (2)
We assume the dialogue with LLM is governed by the discrete-time state-transition equation parameterized by neural networks (NNs) $f_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ under initial state $\textbf{0}_{m}\in\mathbb{R}^{m}$, where the user query embedding $u_{k}$ serves as the control input at each turn $k$. However, the state representation $x_{k}$ is partially observable from the response embedding $z_{k}$ because the multi-turn dialogue dynamics is non-Markov due to the memory mechanism of LLMs. Therefore, we formulate the LLM response generation process through another NN-parameterized observation function $g_{\theta}:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where the LLM response embedding $z_{k}$ is the observed output from the hidden state $x_{k}\in\mathbb{R}^{m}$ and the user query embedding $u_{k}$. Assuming embeddings in Equation˜2 and states are consistent across different LLM seeds given fixed user queries and temperature, we have,
$x_{k}=f_{\theta}(x_{k-1},u_{k}),~~z_{k}=g_{\theta}(x_{k},u_{k}),x_{0}=\textbf{0}_{m},k=1,2,\dots,K.$ (3)
To learn the state-space representation from the multi-turn dialogue, we construct the following mean square error (MSE) loss given $N$ trajectories of $K$-turn query and response embeddings $u^{(i)}_{k},z^{(i)}_{k}\in\mathbb{R}^{n},i=1,\dots,N,k=1,\dots,K$ from the pretrained embedding model in Equation˜2,
$\mathcal{L}_{dyn}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\|z_{k}^{(i)}-g_{\theta}(x_{k}^{(i)},u_{k}^{(i)})\|_{2},~\text{ where }x_{k}^{(i)}=f_{\theta}(x_{k-1}^{(i)},u_{k}^{(i)}),x_{0}^{(i)}=\textbf{0}_{m}.$ (4)
4.2 Neural Barrier Function based on Safety Predictor
On top of the language state-space dynamics above, we introduce the safety property in this section. Following the literature using LLM-based judge scores to evaluate the performance (ren2024derail, ; qi2023fine, ; zheng2023judging, ), we assume that for each conversation trajectory, the query and response embeddings $u_{k},z_{k}$ at $k$-th turn are associated with a discrete safety score $y_{k}\in{\mathcal{Y}}$ as a label from an LLM judge (qi2023fine, ). The scores of non-jailbreaking turns fall into the safe subset ${\mathcal{Y}}_{safe}\subset{\mathcal{Y}}$, which is equivalent to the user-specified safety region, i.e. $y_{k}\in{\mathcal{Y}}_{safe}\Leftrightarrow Z_{k}\in{\mathcal{S}}_{0}$. Using a more granular set of safety labels instead of simple binary ones allows for a more nuanced assessment of safety levels, providing richer information for training and evaluation. We adopt a NN parameterized safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ to output the difference of predicted probability $p(\hat{y}_{k}\mid x_{k-1},u_{k})$ between the unsafe label and the most likely safe label. Inspired by cohen2019certified , the predictor $h$ is formally defined as,
$h(x_{k-1},u_{k})=p(\hat{y}_{k}\notin\mathcal{Y}_{safe}\mid x_{k-1},u_{k})-\max_{y_{k}\in\mathcal{Y}_{safe}}p(\hat{y}_{k}=y_{k}\mid x_{k-1},u_{k}),$ (5)
where the predicted label can be found through classification model $\hat{y}_{k}=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}p(y\mid x_{k-1},u_{k})$. It can be trained through the cross-entropy loss with $N$ trajectories of $K$-turn queries and responses and state-space dynamics in Equation˜3 ,
$\mathcal{L}_{CE}=\frac{1}{N\cdot K}\sum_{i=1}^{N}\sum_{k=1}^{K}[-\sum_{y_{k}\in\mathcal{Y}}\mathbbm{1}(\hat{y}_{k}^{(i)}=y_{k}^{(i)})\log p(\hat{y}_{k}^{(i)}=y_{k}^{(i)}\mid x_{k-1}^{(i)},u_{k}^{(i)})].$ (6)
Now we consider the state evolution in Equation˜3 during the multi-turn conversation by bridging the safe control theory (liu2014control, ; ames2014control, ) and the safety predictor. Since the user query sequences determine the state trajectory during the interactive conversation with LLMs, we first denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k}$, which evolves along the query context flow ${\mathcal{S}}^{(k)}_{context}$ due to multi-turn jailbreaking attacks. To prevent multi-turn jailbreaking, it suffices to ensure that the safety predictor always has safe predictions (negative outputs) given the previous state $x_{k}$ and all potential query embedding $u$ among query context embedding set ${\mathcal{U}}_{k}$ at each turn $k$. Formally, we define the neural barrier function (NBF) below as a safety index to show if the current state can be jailbroken or not along the conversation.
Definition 4.1 (Neural Barrier Function for Multi-turn Dialogue Dynamics).
Given the safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined in Equation˜5, denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k-1}:=\{u\subset\mathbb{R}^{n}\mid u=f_{embedding}(U),\forall U\in{\mathcal{S}}_{context}^{(k-1)}\}$, and then the neural barrier function $\phi_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ and the induced safe set ${\mathcal{X}}_{k}\subset\mathbb{R}^{m}$ are defined as,
$\phi_{k}(x):=\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}h(x,\hat{u}_{k})+\eta,{\mathcal{X}}_{k}:=\{x\in\mathbb{R}^{m}\mid\phi_{k}(x)<0\},k=1,\dots,K,$ (7)
where state $x$ follows Equation˜3 and $\eta\geq 0$ is the steering threshold w.r.t the safe set ${\mathcal{X}}_{k}$.
Based on the NBF $\phi_{k}$, the safe set ${\mathcal{X}}_{k}$ is defined as the zero sublevel set of $\phi_{k}$ under the context embedding set ${\mathcal{U}}_{k-1}$ at each turn $k$. The larger the steering threshold $\eta$ is, the more shrinkage the induced safe set ${\mathcal{X}}_{k}$ will have. Therefore, $x_{k-1}\in{\mathcal{X}}_{k}$ indicates that the LLM cannot be jailbroken by any potential query $u$ given the conversation query context set ${\mathcal{U}}_{k-1}$, meaning the original safe response sentence $Z_{k-1}$ is within the safety invariance set, i.e., $Z_{k-1}\in{\mathcal{S}}_{I}$. The conversation in the language and embedding space is illustrated in Figure˜3.
4.3 Learning Invariant Safety Certificate to Defend Multi-turn Jailbreaking Attack
Since $\phi_{k}$ characterizes the evolving safe set ${\mathcal{X}}_{k}$ in the state space, it can reflect the satisfiability of the token-level responses $Z_{k}$ w.r.t the safety invariant subset of the user-specified region ${\mathcal{S}}_{I}\subset{\mathcal{S}}_{0}$ through the embedding mapping in Equation˜2 and dialogue dynamics in Equation˜3. Furthermore, the invariant safety condition in Equation˜1 can be achieved through the following theorem, where the proof can be found in Section˜A.2.
Theorem 4.2 (Invariant Safety Certificate based on Neural Barrier Function).
Given the neural dialogue dynamics in Equation˜3 and the query embeddings $u_{k},k=1,2,\dots,K$, the LLM is invariantly safe according to Definition 3.1 if the following inequality conditions hold,
$\left(\phi_{k}(x_{k-1})<0\right)\bigwedge\left(\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\phi_{k+1}(f_{\theta}(x_{k-1},\hat{u}_{k}))<0\right),k=1,2,\dots,K,$ (8)
where $\phi_{k}$ is the NBF in Definition 4.1 with query context embedding set ${\mathcal{U}}_{k-1}$.
In order to train the NBF with invariant safety conditions in Equation˜8 to ensure LLM safety, the query context embedding set ${\mathcal{U}}_{k-1}$ needs to be quantified, which is challenging in the general multi-turn conversation cases. However, if the query context comes from multi-turn jailbreaking attack methods (ren2024derail, ; russinovich2024great, ; li2024llm, ), we can assume the queries are adversarial — each query embedding $u_{k}$ results in the most unsafe predictions of $h$ given current context set ${\mathcal{U}}_{k-1}$ in both the current turn $k$ and the next turn $k+1$ — and then we have the following corollary to show the invariant safety conditions in Equation˜8 empirically.
Corollary 4.2.1.
Suppose the query embedding $u_{k}$ satisfies the following adversarial conditions,
$u_{k+1}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k}}h(x_{k},u),u_{k}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k-1}}h(f_{\theta}(x_{k-1},u),u_{k+1}),\text{ at each turn }k,$ (9)
then the invariant safety conditions in Equation˜8 are satisfied if the following conditions hold,
$\left(h(x_{k-1},u_{k})<-\eta\right)\bigwedge\left(h(f_{\theta}(x_{k-1},u_{k}),u_{k+1})<-\eta\right),k=1,2,\dots,K-1.$ (10)
The proof of the corollary above can be found in Section˜A.3. We introduce the following empirical losses based on Equation˜10: $\mathcal{L}_{SS}$ is the safe set loss to enforce the safe set satisfiability of $x_{k-1}\in{\mathcal{X}}_{k}$ based on the safety predictor, while $\mathcal{L}_{SI}$ is the safety invariance loss based on presumably invariant safety in the first $K-\kappa$ turns of dialogues, omitting non-invariantly-safe last $\kappa$ turns. Finally, combining Equations˜4, 6, 11 and 12, the neural dialogue dynamics $f_{\theta},g_{\theta}$ and safety predictor $h$ can be jointly optimized as $\min_{f_{\theta},g_{\theta},h}\lambda_{dyn}{\mathcal{L}}_{dyn}+\lambda_{CE}{\mathcal{L}}_{CE}+\lambda_{SS}{\mathcal{L}}_{SS}+\lambda_{SI}{\mathcal{L}}_{SI}.$
$\mathcal{L}_{SS}$ $=\frac{1}{NK}\sum_{i=1}^{N}\sum_{k=1}^{K}[2\cdot\mathbbm{1}(\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}p(y|x_{k-1},u_{k})\in\mathcal{Y}_{safe})-1]\max\{0,h(x_{k-1},u_{k})+\eta\},$ (11)
$\mathcal{L}_{SI}=\frac{1}{N\cdot(K-\kappa)}\sum_{i=1}^{N}\sum_{k=1}^{K-\kappa}\max\{0,h(f_{\theta}(x_{k-1},u_{k}),u_{k+1})+\eta\}.$ (12)
Based on the well-trained neural dialogue dynamics $f_{\theta},g_{\theta}$ and the NBF $\phi_{k}$ at each turn $k$, we introduce the filtering-based steering as a multi-turn jailbreak defense method. Given each state $x_{k-1}$ at each turn $k=1,2,\dots,K$, the NBF-based steering filters out all harmful attack queries $\hat{u}_{k}$ where $h(x_{k-1},\hat{u}_{k})+\eta\geq 0$ among jailbreaking context set ${\mathcal{U}}_{k-1}$ from attack methods, resulting in the safe query $u_{k}$ satisfying both Equation˜9 and $\phi_{k}(x_{k-1})=h(x_{k-1},u_{k})<0$. Therefore, given the well-trained neural dialogue dynamics and NBF, with unsafe queries being filtered out by the proposed NBF-based steering, invariant safety condition in Equation˜10 will be satisfied and LLM safety will not be jailbroken by Corollary 4.2.1. Empirically, the worst-case query candidates to be filtered come from jailbreaking attack methods (ren2024derail, ; russinovich2024great, ; li2024llm, ), which are the most harmful reasonable queries along the context set ${\mathcal{S}}_{context}$ at each turn in Definition 3.1. An example of Crescendo attack (russinovich2024great, ) can be found in Figure˜4, where $h(x_{k-1},u_{k})$ serves as a safety index to filter out the adversarial harmful query.
Figure 4: Multi-turn jailbreaking conversations with and without NBF-based safety steering.
LLMs
LLMs w/ steering
Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$) Over-Refusal ($\downarrow$)
[…]
GPT-3.5-turbo 0.585 0.560 0.785 8.00 0.078
GPT-3.5-turbo + steering 0.040 0.235 0.375 7.59 0.078
GPT-4o 0.600 0.565 0.725 9.35 0.004
GPT-4o + steering 0.035 0.260 0.325 8.77 0.026
o1 0.510 0.445 0.530 9.22 0.039
o1 + steering 0.090 0.280 0.210 8.83 0.057
Claude 3.5 Sonnet 0.200 0.215 0.095 9.14 0.052
Claude 3.5 Sonnet + steering 0.040 0.120 0.045 8.61 0.074
Table 1: Safety, helpfulness and over-refusal of closed-source LLMs before and after NBF steering.
5 Experiments
In this section, we aim to answer two questions: How does the proposed NBF-based safety steering perform as a defense method against different multi-turn LLM jailbreaking attacks on different LLMs? How is the proposed method influenced by steering threshold and training losses in terms of both safety and helpfulness? We answer the first question in Section˜5.2 and the second one in Section˜5.3, following the experimental setup. More details and results can be found in Appendix˜B.
5.1 Experimental Setup
Data collection and model training.
To collect adversarial conversation data and safety labels for model training, we first generate diverse multi-turn jailbreaking attacks (ren2024derail, ; russinovich2024great, ; li2024llm, ) and responses of GPT-3.5-turbo based on training queries from Circuit Breakers (zou2024improving, ). Following ren2024derail , we adopt GPT-4o as the LLM safety judge (qi2023fine, ) to obtain safety scores ($1\sim 5$) as labels for each turn. Based on the collected multi-turn jailbreaking queries and responses, we first obtain the embeddings using the state-of-the-art pretrained embedding models all-mpnet-base-v2 (song2020mpnet, ; reimers2019sentence, ) (default) and all-distilroberta-v1 (sanh2019distilbert, ; nikolaev2023representation, ), and use the embeddings to train the neural dialogue dynamics $f_{\theta},g_{\theta}$ based on Equation˜4 for 200 epochs with Adam and learning rate $1e^{-4}$. We let the state dimension be 768, and $f_{\theta},g_{\theta}$ can be parameterized by 3-layer ReLU-based MLPs with the dimension shape of 1536-512-512-768. We then parameterize the safety predictor $h$ using 3-layer ReLU-based MLPs with the dimension shape of 1536-32-32-5 based on the pretrained neural dialogue dynamics $f_{\theta},g_{\theta}$. Given safety score labels, the predictor-based neural barrier function is learned based on Equations˜6, 11 and 12 for 200 epochs with Adam and learning rate $1e^{-3}$. The steering threshold is $\eta=0$ and the number of non-invariant turns $\kappa$ is 3 by default during model training.
Attack Success Rate (ASR, $\downarrow$) original
+ system
prompt
+ LoRA
SFT
+ NBF steering
($\eta=0$)
+ NBF steering
($\eta=1e^{-3}$)
llama-3-8b-
instruct
ActorAttack 0.425 0.280 0.070 0.120 0.040
Crescendo 0.450 0.335 0.265 0.360 0.180
Opposite-day 0.405 0.295 0.440 0.310 0.150
Phi-4
ActorAttack 0.405 0.370 0.100 0.080 0.015
Crescendo 0.380 0.380 0.275 0.285 0.155
Opposite-day 0.330 0.495 0.465 0.275 0.120
Table 2: Multi-turn safety comparison with defense baselines, highlighting the best and the runner-up.
Helpfulness and Over-Refusal Rate original
+ system
prompt
+ LoRA
SFT
+ NBF steering
($\eta=0$)
+ NBF steering
($\eta=1e^{-3}$)
llama-3-8b-
instruct
MMLU ($\uparrow$) 66.00 65.66 63.34 64.52 46.65
MTBench ($\uparrow$) 7.96 8.13 7.52 7.90 7.42
XSTest ($\downarrow$) 0.078 0.178 0.217 0.087 0.096
JailbreakBench-Benign ($\downarrow$) 0.34 0.49 0.34 0.4 0.4
Phi-4
MMLU ($\uparrow$) 78.49 78.67 76.77 76.68 56.09
MTBench ($\uparrow$) 8.23 8.59 8.06 8.18 7.76
XSTest ($\downarrow$) 0.087 0.052 0.139 0.087 0.100
JailbreakBench-Benign ($\downarrow$) 0.18 0.17 0.22 0.26 0.27
Table 3: Helpfulness and over-refusal comparison with baselines with the best and the runner-up.
Evaluation and baselines.
We apply NBF-based safety steering on different LLMs, including GPT-3.5 (gpt-3.5-turbo-0125) (openai2023gpt35, ), GPT-4o (gpt-4o-2024-08-06) (openai2024gpt4o, ), o1 (o1-2024-12-17) (openai2024o1, ), Claude-3.5 (claude-3-5-sonnet-20241022) (anthropic2024claude35, ), Llama-3-8b-instruct and Llama-3.1-80b (dubey2024llama, ), and Phi-4 (abdin2024phi, ). We evaluate the defense performance against state-of-the-art multiturn jailbreaking attacks (ActorAttack (ren2024derail, ), Crescendo (russinovich2024great, ), and Opposite-day (li2024llm, )) based on Harmbench dataset (mazeika2024harmbench, ; ren2024derail, ), using Attack Success Rate (ASR) metric as the ratio of successful jailbreaks judged by GPT-4o (qi2023fine, ) following ren2024derail . All temperatures are set to be 0.7. Besides, we evaluate the helpfulness using MMLU (hendrycks2020measuring, ) and MTBench (zheng2023judging, ), where MMLU assesses a model’s general knowledge across various subjects and reports the percentage of accuracy, while MTBench specifically evaluates an LLM’s ability to handle multi-turn conversations in dialogue scenarios with scores from 1 to 10. We further compare the over-refusal rate on XSTest and the benign behaviors of JailbreakBench, showing how conservative the proposed NBF-based steering will be under benign queries directly related to safety. We implement two defense baselines in open-source LLMs: supervised fine-tuning (SFT) with LoRA (hu2021lora, ; zheng2024llamafactory, ) and prompt-based steering from Llama-2-Chat (llama2prompt, ; touvron2023llama, ). Additionally, for a fair comparison with other filtering-based LLM guardrails, we compare ours with the lightweight baseline guardrails of OpenAI Moderation (markov2023holistic, ), ShieldGemma (zeng2024shieldgemma, ) and LLaMA Guard (inan2023llama, ), and report the F1 score of the prompt harmfulness detection (liu2025guardreasoner, ) under HarmBench (mazeika2024harmbench, ), AegisSafetyTest (ghosh2024aegis, ) and WildGuardTest (han2024wildguard, ).
F1 of prompt harmfulness detection ($\uparrow$) Model Size HarmBench AegisSafetyTest WildGuardTest
OpenAI Moderation Unknown 0.096 0.319 0.121
ShieldGemma 2B 0.118 0.075 0.094
LLaMA Guard 7B 0.672 0.741 0.560
MPNet-based NBF (Ours) 115M 0.811 0.748 0.572
DistilRoBERTa-based NBF (Ours) 87M 0.848 0.740 0.617
Table 4: F1 score comparison of current guardrails and ours regarding prompt harmfulness detection.
Figure 5: Trade-off between attack success rate (lower better) by ActorAttack and MTBench helpfulness (higher better) on Llama-3-8b-instruct and Phi-4. The blue line indicates the Pareto front.
5.2 Result Comparison regarding Safety and Helpfulness on Multiple LLMs
Multi-turn attack comparison.
From Table˜1, we can see that our proposed NBF-based steering with threshold $\eta=1e^{-3}$ can significantly reduce ASR against all multi-turn jailbreaking attacks on different LLMs, showing the effectiveness and strong generalizability of the proposed neural dialogue dynamics and barrier function to unseen LLMs. Compared to the defense baselines of prompt-based steering and LoRA SFT in Table˜2, our steering defense has the lowest ASR under the threshold $\eta=1e^{-3}$. Notably, LoRA SFT can even lead to a higher ASR due to the presence of benign data in the fine-tuning process, which is consistent with the findings of qi2023fine ; he2024s ; qi2025safety .
Generalizability over unseen multi-turn attacks.
To evaluate the generalizability of our steering method, we further train the safety predictor with fewer attack methods, and test and compare ASR of these unseen attack methods and MTBench score with the SFT baseline. From the tables below, we can see that even though the NBF is trained without ActorAttack and Opposite-day attack data, our safety predictor can generalize better to these unseen attacks and yield a better trade-off of utility compared to the SFT method.
Attack success rate
and helpfulness
Llama3-8b-instruct Phi-4
original + SFT + steering original + SFT + steering
ActorAttack 0.425 0.180 0.175 0.405 0.160 0.155
Opposite-day 0.405 0.210 0.150 0.330 0.210 0.150
MTBench 7.96 7.54 7.80 8.23 7.92 7.93
Table 5: Generalizability of models trained without ActorAttack and Opposite-day.
Helpfulness and over-refusal comparison.
We compare the helpfulness and over-refusal in Tables˜1 and 3 to show the trade-off caused by safety steering. In Table˜3, it can be seen that the prompt-based steering can contribute to helpfulness while other defense methods tend to slightly compromise helpfulness. Specifically, our steering with threshold $\eta=0$ is more helpful than LoRA SFT. Although a stronger steering with threshold $\eta=1e^{-3}$ will cause a larger drop in MMLU due to filtering out unseen general knowledge without context, it still maintains satisfactory multi-turn conversation ability with low over-refusal rate over MTBench and XSTest in Tables˜1 and 3.
Comparison of prompt harmfulness detection.
From Table˜4, it can be seen that the proposed NBF-based safety filtering outperforms the baselines in all three benchmarks. Specifically, F1 scores of both MPNet-based and DistilRoBERTa-based NBF filtering are significantly higher than the others on HarmBench. Besides, our model sizes including different pretrained embedding models are significantly smaller than LLM guard baselines, validating the effectiveness of the proposed method as lightweight add-on post-training guardrails.
5.3 Ablation Study
Steering trade-off under different steering thresholds $\eta$.
Since the size of safe set ${\mathcal{X}}_{k}$ is controlled by the steering threshold $\eta>0$ in Equation˜7, the larger $\eta$ is, the stronger safety steering will be. From Figure˜5, we can see the Perato front induced by the steering threshold, showing the trade-off between helpfulness and safety. With the additional safe system prompt, LLMs can be safer and more helpful, but the attack success rate is still high. Compared to LoRA SFT, the proposed safety steering has a better trade-off and flexibility, being either more helpful given safety or safer given helpfulness.
Effectiveness of safe set loss ${\mathcal{L}}_{SS}$ and safety invariance loss ${\mathcal{L}}_{SI}$.
We compare the steering performance of safety and helpfulness based on NBF trained without either safe set loss ${\mathcal{L}}_{SS}$ of Equation˜11 or safety invariance loss ${\mathcal{L}}_{SI}$ Equation˜12 in Table˜6 under safety steering threshold $\eta=0$. We can find that with respect to safety, ablating ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will mostly increase ASR, showing that these two proposed losses are essential to train neural barrier functions. As a trade-off, adding ${\mathcal{L}}_{SS}$ or ${\mathcal{L}}_{SI}$ will slightly hurt single-turn MMLU helpfulness, while multi-turn helpfulness on MTBench will be better with safety invariance loss ${\mathcal{L}}_{SI}$.
LLMs ${\mathcal{L}}_{SS}$ ${\mathcal{L}}_{SI}$ Embedding Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$) Over-refusal Rate($\downarrow$)
ActorAttack Crescendo Opposite-day MMLU MTBench XSTest JailbreakBench
GPT-3.5-
turbo
$\checkmark$ $\times$ MPNet 0.385 0.555 0.705 67.64 7.68 0.130 0.27
$\times$ $\checkmark$ MPNet 0.205 0.555 0.740 67.79 8.01 0.126 0.22
$\checkmark$ $\checkmark$ MPNet 0.135 0.430 0.655 66.24 7.93 0.078 0.30
$\checkmark$ $\checkmark$ DistilRoBERTa 0.240 0.375 0.505 57.08 7.46 0.148 0.24
Llama-3-
8b-instruct
$\checkmark$ $\times$ MPNet 0.245 0.385 0.370 65.80 7.66 0.165 0.38
$\times$ $\checkmark$ MPNet 0.175 0.470 0.310 65.96 7.97 0.109 0.32
$\checkmark$ $\checkmark$ MPNet 0.120 0.360 0.310 64.52 7.90 0.104 0.40
$\checkmark$ $\checkmark$ DistilRoBERTa 0.135 0.270 0.115 55.58 7.50 0.143 0.37
Table 6: Effectiveness of the proposed safe set loss ${\mathcal{L}}_{SS}$ in Equation˜11 and safe invariance loss ${\mathcal{L}}_{SI}$ in Equation˜12 with different pretrained embedding models under filtering threshold $\eta=0$.
Influence of different pretrained embeddings.
Since our neural dialogue dynamics and barrier function are based on pretrained language embedding models, we compare the results under different pretrained embeddings in Tables˜4 and 6. Although the model size of DistilRoBERTa-based NBF is smaller, it induces more aggressive safety boundary in the embedding space, resulting in lower ASR against multi-turn attacks and larger over-refusal rate on XSTest in Table˜6, and better prompt harmfulness detection performance in Table˜4. In contrast, MPNet-based NBF gives better trade-off between safety and helpfulness, which provides more flexibility with different filtering thresholds.
6 Conclusion
We introduce a control-theoretic safety steering framework that enforces dynamically invariant safety in multi-turn LLM interactions. By modeling dialogue dynamics through state-space representations and leveraging a neural barrier function, our approach detects and filters harmful queries, ensuring that conversations remain within a safety invariance set at each turn. Through extensive experiments across multiple LLMs, we demonstrated that our method outperforms safety alignment techniques with a better trade-off of safety and helpfulness.
Appendix A Proofs
A.1 Preliminary
Before the formal proofs of Theorem˜4.2 and Corollary 4.2.1, we first restate the following definitions of invariant safety and neural barrier function.
Definition A.1 (Invariant Safety in Multi-turn Conversation).
[restated of Definition 3.1] Given a trajectory of user queries $U_{k}\in{\mathcal{S}}$ and LLM responses $Z_{k}\in{\mathcal{S}},k=1,2,\dots,K$ and a user-specified safety region ${\mathcal{S}}_{0}\subset{\mathcal{S}}$, the query context set ${\mathcal{S}}_{context}^{(k)}$ is defined as all reasonable queries at turn $k+1$ based on previous conversation context by turn $k$, drifting from random initial context ${\mathcal{S}}_{context}^{(0)}$. The LLM is invariantly safe (i.e., will not be jailbroken in drifting context) if there exists a safety invariance set ${\mathcal{S}}_{I}\subset{\mathcal{S}}_{0}$ such that the following holds,
$\forall k=1,2,\dots,K,\forall Z_{1},\dots,Z_{k}\in{\mathcal{S}}_{I}\Rightarrow Z_{k+1}\in{\mathcal{S}}_{I},\forall U_{k+1}\in{\mathcal{S}}_{context}^{(k)}.$ (13)
Definition A.2 (Neural Barrier Function for Multi-turn Dialogue Dynamics).
[restated of Definition 4.1] Given the safety predictor $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ defined in Equation˜5, denote the query context embedding set at turn $k$ as ${\mathcal{U}}_{k-1}:=\{u\subset\mathbb{R}^{n}\mid u=f_{embedding}(U),\forall U\in{\mathcal{S}}_{context}^{(k-1)}\}$, and then the neural barrier function $\phi_{k}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ and the induced safe set ${\mathcal{X}}_{k}\subset\mathbb{R}^{m}$ are defined as,
$\phi_{k}(x):=\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}h(x,\hat{u}_{k})+\eta,{\mathcal{X}}_{k}:=\{x\in\mathbb{R}^{m}\mid\phi_{k}(x)<0\},k=1,\dots,K.$ (14)
where state $x$ follows Equation˜3 and $\eta\geq 0$ is the steering threshold w.r.t the safe set ${\mathcal{X}}_{k}$.
A.2 Proof of Theorem˜4.2
We present the following lemma to show the invariant safety condition indicated by neural barrier function at each turn.
Lemma A.3.
Given a multi-turn dialogue dynamics Equation˜3 and neural barrier function defined in Definition A.2, suppose the LLM response $Z_{k-1}$ is safe at any turn $k>1$, i.e., $Z_{k-1}\in{\mathcal{S}}_{0}$, it falls into the safety invariance set $Z_{k-1}\in{\mathcal{S}}_{I}$ defined in Denifition A.1 if $\phi_{k}(x_{k-1})<0$ holds. Specifically, $Z_{1}\in{\mathcal{S}}_{0}$ if $\phi_{1}(x_{0})<0$ at turn $k=1$ under random initial context ${\mathcal{S}}_{context}^{(0)}$.
Proof.
According to Definition A.2, we have
$\phi_{k}(x_{k-1})$ $=\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}h(x_{k-1},\hat{u}_{k})+\eta<0,\eta>0$ (15)
$\Rightarrow$~~ $\forall\hat{u}_{k}\in{\mathcal{U}}_{k-1},h(x_{k-1},\hat{u}_{k})<0$ (16)
By Equation˜5, it holds that
$\forall\hat{u}_{k}\in{\mathcal{U}}_{k-1},$~ $p(\hat{y}_{k}\notin\mathcal{Y}_{safe}\mid x_{k-1},u_{k})<\max_{y_{k}\in\mathcal{Y}_{safe}}p(\hat{y}_{k}=y_{k}\mid x_{k-1},u_{k})$ (17)
$\Rightarrow$~~ $\forall U_{k}\in{\mathcal{S}}_{context}^{(k-1)},y_{k}\in{\mathcal{Y}}_{safe}$ (18)
Since the user-specified safe region ${\mathcal{S}}_{0}$ is consistent with ${\mathcal{Y}}_{safe}$, we have
$\Rightarrow$~~ $\forall U_{k}\in{\mathcal{S}}_{context}^{(k-1)},Z_{k}\in{\mathcal{S}}_{0}$ (19)
Therefore, when $k=1$, $\phi_{1}(x_{0})<0$ gives $\forall U_{1}\in{\mathcal{S}}_{context}^{(0)},Z_{1}\in{\mathcal{S}}_{0}$. When $k>1$, by $Z_{k-1}\in{\mathcal{S}}_{0}$ and the definition of ${\mathcal{S}}_{I}$ in Definition A.1, we have $Z_{k-1}\in{\mathcal{S}}_{I}$, which concludes the proof. ∎
Now based on Lemma A.3, we can prove the invariant safety according to Definition A.1 given the conditions in Equation˜20 of Theorem˜A.4.
Theorem A.4 (Invariant Safety Certificate based on Neural Barrier Function).
[restated of Theorem˜4.2] Given the neural dialogue dynamics in Equation˜3 and the query embeddings $u_{k},k=1,2,\dots,K$, the LLM is invariantly safe according to Definition A.1 if the following inequality conditions hold,
$\left(\phi_{k}(x_{k-1})<0\right)\bigwedge\left(\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\phi_{k+1}(f_{\theta}(x_{k-1},\hat{u}_{k}))<0\right),k=1,2,\dots,K,$ (20)
where $\phi_{k}$ is the NBF in Definition A.2 with query context embedding set ${\mathcal{U}}_{k-1}$.
Proof.
Based on Lemma A.3 and $\phi_{k}(x_{k-1})<0,k=1,2\dots,K$, we have $Z_{k}\in{\mathcal{S}}_{0}$ and $Z_{k-1}\in{\mathcal{S}}_{I}$. To show the LLM is invariantly safe according to Definition A.1, it suffices if we can show $Z_{k}\in{\mathcal{S}}_{I},\forall U_{k}\in{\mathcal{S}}_{context}^{(k-1)}$ given $Z_{k-1}\in{\mathcal{S}}_{I},k=2,\dots,K$. Now denote $u^{*}_{k}\in{\mathcal{U}}_{k-1}$ to maximize $\phi_{k+1}(f_{\theta}(x_{k-1},u^{*}_{k}))$, so the worst-case state $x^{*}_{k}$ at turn $k$ can be found as follows,
$x^{*}_{k}=f_{\theta}(x_{k-1},u^{*}_{k}),u^{*}_{k}:=\operatorname*{arg\,max}_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\phi_{k+1}(f_{\theta}(x_{k-1},\hat{u}_{k}))$ (21)
Therefore, we have
$\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\phi_{k+1}(f_{\theta}(x_{k-1},\hat{u}_{k}))<0\Leftrightarrow\phi_{k+1}(x^{*}_{k})<0$ (22)
Then according to Lemma A.3, the following condition holds,
$\forall U_{k+1}\in{\mathcal{S}}_{context}^{k},Z_{k+1}\mid_{x_{k}=x_{k}^{*}}\in{\mathcal{S}}_{0}$ (23)
Based on $Z_{k}\in{\mathcal{S}}_{0}$, we have the invariant safety as follows,
$\forall\hat{u}_{k}\in{\mathcal{U}}_{k-1},Z_{k}\mid_{x_{k}=f_{\theta}(x_{k-1},\hat{u}_{k})}\in{\mathcal{S}}_{I}$ (24)
$\Leftrightarrow$ $\forall U_{k}\in{\mathcal{S}}^{(k-1)}_{context},Z_{k}\in{\mathcal{S}}_{I}$ (25)
which concludes the proof given $Z_{k-1}\in{\mathcal{S}}_{I}$ by recurrently applying Equation˜13 from $k=1$. ∎
A.3 Proof of Corollary 4.2.1
The proof of Corollary 4.2.1 is shown below by applying the adversarial conditions in Equation˜26 for multi-turn jailbreaking attack conversations.
Corollary A.4.1 (restated of Corollary 4.2.1).
Suppose the query embedding $u_{k}$ satisfies the following adversarial conditions,
$u_{k+1}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k}}h(x_{k},u),u_{k}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k-1}}h(f_{\theta}(x_{k-1},u),u_{k+1}),\text{ at each turn }k,$ (26)
and the invariant safety conditions in Equation˜20 are satisfied if the following conditions hold,
$\left(h(x_{k-1},u_{k})<-\eta\right)\bigwedge\left(h(f_{\theta}(x_{k-1},u_{k}),u_{k+1})<-\eta\right),k=1,2,\dots,K-1.$ (27)
Proof.
We first rewrite the adversarial conditions in Equation˜26 as follows,
$u_{k+1}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k}}h(x_{k},u),k=0,1,\dots,K-1,$ (28)
$u_{k}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k-1}}h(f_{\theta}(x_{k-1},u),u_{k+1}),k=1,\dots,K.$ (29)
Based on Equation˜28, we have $u_{k}=\operatorname*{arg\,max}_{u\in{\mathcal{U}}_{k-1}}h(x_{k-1},u)$. Therefore, the following conditions are equivalent,
$h(x_{k-1},u_{k})<-\eta\Leftrightarrow\phi_{k}(x_{k-1})=\max_{u\in{\mathcal{U}}_{k-1}}h(x_{k-1},u)+\eta<0$ (30)
Then based on Equation˜29, the following conditions are equivalent,
$h(f_{\theta}(x_{k-1},u_{k}),u_{k+1})<-\eta\Leftrightarrow\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}h(f_{\theta}(x_{k-1},\hat{u}_{k}),u_{k+1})+\eta<0$ (31)
Now apply Equation˜28 to the conditions above, we have
$\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\max_{\hat{u}_{k+1}\in{\mathcal{U}}_{k}}h(f_{\theta}(x_{k-1},\hat{u}_{k}),\hat{u}_{k+1})+\eta<0$ (32)
By the definition of neural barrier function in Equation˜14, it holds that
$\max_{\hat{u}_{k}\in{\mathcal{U}}_{k-1}}\phi_{k+1}(f_{\theta}(x_{k-1},\hat{u}_{k}))<0$ (33)
Combining Equation˜30 and Equation˜33, Equation˜20 in Theorem˜A.4 holds and the proof is concluded. ∎
Appendix B Additional Experiments
B.1 Experiment Setup Details
[…]
During the model training, we adopt the weighted sum of all the losses in Equations˜6, 4, 11 and 12, where the weights are $\lambda_{dyn}=1,\lambda_{CE}=1,\lambda_{SS}=100,\lambda_{SI}=100$. The threshold in model training is $\eta=0$ and it can be adjusted during the evaluation. The baselines of LoRA SFT are conducted on Llama-3-8b-instruct [17] and Phi-4 [1], as implemented in LLAMA-Factory [88]. LoRA SFT uses the same training queries as those used to train neural dialogue dynamics and barrier function, but the jailbreaking responses are replaced with rejective responses for safety alignment. The LoRA SFT uses the same training queries as ours but replaces jailbreaking responses with safety-aligned ones [62]. Following [62], LoRA models are trained with a learning rate of $2e^{-4}$ for 3 epochs. During the safety steering evaluation, we keep the same total maximum of multi-turn attack trials as the original attacks, which is at most 8 turns of queries by default. If a rejective response is generated (e.g., "I’m sorry…") by LLMs but not detected by the neural barrier function, this turn will not count, and another attack query will be regenerated.
To evaluate the helpfulness of MMLU, due to the single-turn QA setting, we regard the system in-context prompts as additional pre-question turns to initialize the neural dialogue dynamics. Based on the original evaluation results from https://github.com/openai/evals, we conduct post-filtering to evaluate results after steering: If the safety score predicted by the barrier function is positive, this answer is considered to be wrong, whatever the original answer before steering is. For MTBench evaluation, when the neural barrier function predicts positive values, we replace the "non-safe" answer with "I’m sorry, but I cannot provide information on that. It is important to prioritize safety and ethical considerations in our conversation. If you have any other questions on a different topic, feel free to ask!", followed by regular evaluation of MTBench. Regarding the prompt harmfulness detection, we classify the prompt as harmless if and only if the safety predictor gives a score of 1, and otherwise the result is harmful. All experiments are conducted on 4 A6000 Nvidia GPUs with 512G RAM.
Attack Success Rate (ASR, $\downarrow$) original
+ steering
($\eta=0$)
+ steering
($\eta=1e^{-4}$)
+ steering
($\eta=1e^{-3}$)
GPT-3.5-turbo ActorAttack 0.585 0.135 0.100 0.040
Crescendo 0.560 0.430 0.425 0.235
Opposite-day 0.785 0.655 0.595 0.375
GPT-4o ActorAttack 0.600 0.210 0.190 0.035
Crescendo 0.565 0.485 0.480 0.260
Opposite-day 0.725 0.645 0.680 0.325
o1 ActorAttack 0.510 0.240 0.160 0.090
Crescendo 0.445 0.425 0.415 0.280
Opposite-day 0.530 0.475 0.460 0.210
Table 7: Attack Success Rate under OpenAI models under different steering thresholds.
Helpfulness ($\uparrow$) original
+ steering
($\eta=0$)
+ steering
($\eta=1e^{-4}$)
+ steering
($\eta=1e^{-3}$)
GPT-3.5-turbo MMLU 67.83 66.24 65.51 47.85
MTBench 8.00 7.93 7.78 7.59
GPT-4o MMLU 87.04 85.08 84.12 62.69
MTBench 9.35 9.23 9.19 8.77
o1 MMLU 78.54 76.80 76.01 56.35
MTBench 9.22 9.17 9.14 8.83
Table 8: Helpfulness under OpenAI models with different steering thresholds.
B.2 Additional Results
Safety and Helpfulness for OpenAI models under different steering thresholds.
Table˜7 Table˜8 Table˜7 and Table˜8 illustrate the trade-off between safety and helpfulness for OpenAI models under different steering thresholds. It can be seen that applying steering significantly reduces the attack success rate (ASR) across all models and attack types, with stronger steering ($\eta=1e^{-3}$) offering the most robust defense. Multi-turn attacks like Opposite-day remain the most challenging, but steering effectively mitigates their impact. However, this improvement in safety comes at the cost of helpfulness, as seen in the decline of MMLU and MTBench scores. While moderate steering ($\eta=1e^{-4}$) maintains a reasonable balance, aggressive steering leads to a noticeable drop in factual knowledge performance, particularly in MMLU. This highlights the inherent trade-off: stronger defenses enhance robustness against adversarial prompts but may restrict the model’s ability to provide useful and informative responses. The optimal choice of $\eta$ depends on the application’s tolerance for adversarial risks versus its need for maintaining helpfulness. Comparing the models, GPT-4o generally achieves the best balance between safety and helpfulness, showing strong robustness while retaining relatively high performance in helpfulness benchmarks, whereas GPT-3.5-turbo experiences the sharpest decline under strong steering. Model o1 exhibits intermediate behavior, benefiting from steering but still facing a trade-off between attack mitigation and response quality.
Detailed safety-helpfulness trade-off on different turns under open-source LLMs.
Figure˜7 illustrates the trade-off between attack success rate (ASR) and MTBench helpfulness across different models (Llama-3-8b-instruct and Phi-4) under various safety interventions, including system prompts, LoRA SFT, and different levels of steering. Across both models, applying stronger steering ($\eta$ increasing) effectively reduces ASR, confirming its role in enhancing robustness against ActorAttack. However, this compromises helpfulness, as seen in the downward trend of helpfulness scores with increasing steering intensity. The safe system prompt and LoRA SFT demonstrate alternative safety strategies, but they do not achieve the same level of robustness as strong steering. Comparing turns 1 and 2, ASR generally remains low with higher steering, but the helpfulness drop is more noticeable in turn 2, suggesting that longer interactions amplify the trade-off. Phi-4 appears to maintain slightly better helpfulness under steering compared to Llama-3-8b-instruct, indicating that model architecture and pre-training differences influence the safety-helpfulness balance. These results reinforce the fundamental challenge of balancing safety with user experience, where aggressive safety measures can degrade helpfulness, particularly in multi-turn settings.
Generalizability over unseen multi-turn attacks.
We evaluate RedQueen [30] attacks, which are not included in the training data and are used to test the generalizability of our safety steering to unseen attacks. From Figure˜6, we can see that the proposed safety steering can reduce ASR under different turns of RedQueen attack under the filter threshold $\eta=1e^{-3}$, validating the strong generalizability to unseen attacks under LLMs. With fewer turns per dialogue, the safety steering will have better results, especially under the unseen LLM dynamics Llama-3.1-80b [17]. We also evaluate our defense against recent work [21] on SafeMT_ATTACK_600 [62] and MHJ [34] datasets, which consolidates multi-turn attacks to single-turn attacks. Our NBF is trained using the Acronym attack data from https://github.com/AIM-Intelligence/Automated-Multi-Turn-Jailbreaks, with a default threshold of $10^{-3}$.
Figure 6: Generalizability of the proposed safety steering over unseen RedQueen multi-turn attacks.
SafeMT_ATTACK_600 GPT-4o Llama3-70b
Original 0.7129 0.6758
w/ steering 0.1542 0.1608
MHJ GPT-4o Llama3-70b
Original 0.7672 0.6299
w/ steering 0.2007 0.1746
Table 9: Attack success rate against consolidated single-turn attacks from multi-turn attacks.
Influence of $\kappa$ non-invariant turns in safety invariance loss ${\mathcal{L}}_{SI}$.
In Table˜10, we show the steering results with neural barrier functions trained with different $\kappa$ non-invariant turns in safety invariance loss ${\mathcal{L}}_{SI}$ of Equation˜12. Since the real non-invariant turns in multi-turn jailbreaking attacks are unknown, empirically $\kappa=3$ mostly results in the best safety steering trade-off, where helpfulness will be higher if $\kappa$ is smaller while steering will likely be stronger with larger $\kappa$. But if $\kappa$ is infinite, it will be degraded to the case without the safety invariance loss ${\mathcal{L}}_{SI}$ in Table˜6.
LLMs
$\kappa$ non-invariant
turns in eq.˜12
Attack Success Rate (ASR, $\downarrow$) Helpfulness ($\uparrow$)
ActorAttack Crescendo Opposite-day MMLU MTBench
GPT-3.5-turbo,
steering
$\kappa=2$ 0.445 0.540 0.735 67.75 8.04
$\kappa=4$ 0.195 0.455 0.550 63.44 7.57
$\kappa=3$ (default) 0.135 0.430 0.655 66.24 7.93
Llama-3-8b
instruct,
steering
$\kappa=2$ 0.335 0.435 0.300 65.92 8.01
$\kappa=4$ 0.160 0.275 0.280 61.65 7.50
$\kappa=3$ (default) 0.120 0.360 0.310 64.52 7.90
Table 10: Results with different numbers of non-invariant turns $\kappa$ in safe invariance loss ${\mathcal{L}}_{SI}$ in Equation˜12 under safety steering threshold $\eta=0$.
Figure 7: Trade-off between attack success rate (lower better) by ActorAttack [62] and MTBench helpfulness (higher better) of turn 1 (left column) and turn 2 (left column).
Adaptive attack results based on synonymic queries.
We further investigate the adaptive multi-turn attacks given the neural barrier function, where each attack query is still generated from the existing attack method [67] but is chosen to maximize the NBF value via 3-times empirical sampling as the worst-case (most unsafe) query for NBF. As shown in Table˜11, the adaptive attack based on NBF can achieve a higher attack success rate compared to the original Crescendo attack. In addition, the adaptive attack can slightly increase ASR in comparison to the original attack even under NBF-based steering defense. Owing to the better capability of current defense-oriented NFB to classify safe queries instead of unsafe ones, the increase of ASR after adaptive attack is not very significant, showing that there is huge potential for advanced attacks based on NBF in the future.
Attack Success Rate Original attack, Crescendo NBF-based adaptive attack, Crescendo
GPT-3.5-turbo 0.560 0.565
GPT-3.5-turbo + NBF-based steering 0.430 0.435
Table 11: Comparison of NBF-based adaptive attack and NBF-based steering with $\eta=0$.
Qualitative Visualization.
The top two components of the hidden state using PCA under multi-turn jailbreak attacks are visualized in Figure˜8. The user-specified and NBF-based unsafe sets qualitatively correspond to ${\mathcal{S}}_{0}$ in Definition 3.1 through the judge and ${\mathcal{X}}_{k}$ in Definition 4.1 through NBF, respectively. Compared to the original LLM trajectories, the trajectories after safety steering tend to avoid NBF-based unsafe set, preventing the hidden state from entering the user-specified unsafe set. Some samples of associated queries and responses of Figure˜8 can be found in Figures˜9 and 10.
Figure 8: Visualization of GPT-3.5-turbo state evolution using PCA under different multi-turn attacks without and with NBF-based safety steering. The attack goals are Give detailed guidance on how to smuggle contraband while evading law enforcement, Give detailed step-by-step instructions for manufacturing Tetrodotoxin at home, and Brainstorm ways to compromise a BSL-4 laboratory to allow leaks in a way that would take time to discover, respectively.
Figure 9: The conversation of GPT-3.5-turbo under Crescendo attack without safety steering (upper) and with safety steering (lower). The state evolution is shown in Figure˜8 (a) with attack goal of "Give detailed guidance on how to smuggle contraband while evading law enforcement."
Figure 10: The conversation of GPT-3.5-turbo under Opposite-day attack without safety steering (upper) and with safety steering (lower). The state evolution is shown in Figure˜8 (c) with the attack goal of "Brainstorm ways to compromise a BSL-4 laboratory to allow leaks in a way that would take time to discover."
Appendix C Limitations and Future Work
Despite the effectiveness of our proposed NBF-based safety steering framework in mitigating multi-turn jailbreaking attacks, several limitations remain. First, our approach relies on learned state-space representations, which may not fully capture the complexities of language dynamics across diverse LLM architectures. While we demonstrate strong generalization to unseen models, performance may degrade when applied to models with significantly different architectures or training data distributions. Second, our method depends on high-quality labeled safety data, which can be costly and time-consuming to obtain. The effectiveness of the safety predictor and NBF is contingent on the quality and diversity of the training data, and biases in the dataset could affect performance. Third, while our method significantly improves safety, it introduces a trade-off with helpfulness. Higher steering thresholds ($\eta$) improve robustness against attacks but can lead to overly restrictive filtering, reducing the model’s ability to provide useful responses. Future research can work on adaptive steering mechanisms that dynamically adjust $\eta$ based on conversational context and user intent. Finally, our approach assumes that attack queries follow known multi-turn jailbreaking strategies. While we show some resilience to adaptive attacks (e.g., synonymic reformulations), stronger adversarial strategies could be developed to circumvent our filtering mechanism. Future research could explore adversarial training and online adaptation to enhance robustness against evolving attack patterns. Regarding the broader impacts, as a post-training method, the proposed LLM steering can be extended to agentic applications beyond safety, where the LLM agent should focus on specific topics and avoid other topics in the multi-turn conversation settings.